Solve for all 0 degrees less than or equal to x less than 360 degrees.
3tan x- sqrt( 3 )=0
2sin x cos x=2sin x
cos^2 x-sin^2 x + 3cos x - 1=0

Found 2 solutions by lwsshak3, ikleyn:
Answer by lwsshak3(11628)

Solve for all 0 degrees less than or equal to x less than 360 degrees.
3tan x- sqrt( 3 )=0
tanx=√3/3
x=30˚, 150˚
..
2sin x cos x=2sin x
cosx=1
x=0
..
cos^2 x-sin^2 x + 3cos x - 1=0
cos^2x-(1-cos^2x)+3cosx-1=0
cos^2x-1+cos^2x+3cosx-1=0
2cos^2x+3cosx-2=0
(2cosx-1)(cosx+2)=0
2cosx-1=0
cosx=1/2
x=60˚, 300˚
..
cosx+2=0
cosx=-2
no solution (-1 ≤ cosx ≤ 1)

Answer by ikleyn(53996)

.
Solve for all 0 degrees less than or equal to x less than 360 degrees.
(a) 3tan x- sqrt( 3 )=0
(b) 2sin x cos x=2sin x
(c) cos^2 x-sin^2 x + 3cos x - 1=0
~~~~~~~~~~~~~~~~~~~~~~


In the post by @lwsshar3, the solution to (a) is incorrect: 150˚ is WRONG answer.
The solution to (b) is incorrect, too: some roots are missed.
I came to bring correct solutions to (a) and (b).


- - - (a) - - -

3tan x- sqrt( 3 )= 0
tanx = √3/3
x = 30˚ or 210˚.

ANSWER to (a). x = 30˚ or 210˚.


- - - (b) - - -

2sin(x)*cos(x) = 2sin(x)
2sin(x)*cos(x) - 2sin(x) = 0
2sin(x)*(cos(x)-1)) = 0

case 1. sin(x) = 0 ---> x = 0˚ or x = 180˚.

case 2. cos(x) = 1 ---> x = 0˚.


ANSWER to (b). x = 0˚ or x = 180˚.